OpenStudy (anonymous):

The geometric sequence that tells the perimeter of each stage in the building of Koch's snowflake has a first term of 3 and a ratio of 4/3 . At the 16th stage in the construction of this snowflake, the length of each tiny side would be about the size of an atom! Using the explicit formula for the nth term of a geometric sequence, what is the length of the perimeter a16 to the nearest whole number?

OpenStudy (dumbcow):

= 3*(4/3)^15 =4^15/3^14 =224

OpenStudy (anonymous):

the snowflake is built on the cantor set if that helps you at all. In fact I think it is a manipulation of the canter founction